# On RoCK blocks of double covers of symmetric and alternating groups and the refined Broué conjecture

### v1

@@ 5 On spin blocks of $\widetilde{S}_{n}$ and $\widetilde{A}_{n}$
Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$. Then $R_{p}Gf$ and $R_{p}Nf$ are Morita superequivalent. More precisely, there is an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$, together with its dual $(X^{R_{p}})^{*}$, inducing a Morita superequivalence between $R_{p}Gb$ and $R_{p}N^{\rho,d}f$ and fulfilling $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$. The bimodule $X^{\mathbb{F}_{p}}$ has endopermutation source.
Proof. By Propositions 3.5 and 3.4, it suffices to show that there exists an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$. By Remark (5.14) we have an isomorphism of $R_{p}$-superalgebras
$R_{p}N^{\rho,d}f\cong B^{\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}}).$
We let $M^{R_{p}}$ be the $(B^{\varnothing,1}_{R_{p}},B^{\varnothing,1}_{R_{p}})$-bimodule
$M^{R_{p}}:=\Omega_{B^{\varnothing,1}_{R_{p}}\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}})^{\rm sop}}^{\ell}(B^{\varnothing,1}_{R_{p}}).$
@@
$\begin{split}M_{N^{\rho,d}}^{R_{p}}&:=B^{\rho,0}_{R_{p}}\boxtimes_{R_{p}}(M^{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})\\ &=B^{\rho,0}_{R_{p}}\boxtimes_{R_{p}}\Big({\rm Ind}_{(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}},B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})_{S_{d}}}^{(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})\otimes_{R_{p}}(B^{\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}_{d}^{R_{p}})^{\rm sop}}(M^{R_{p}})^{\boxtimes d}_{S_{d}}\Big).\end{split}$
Then we have $M_{N^{\rho,d}}^{R}\cong R\otimes_{R_{p}}M_{N^{\rho,d}}^{R_{p}}$ as $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodules. Now by Proposition 3.9, there is an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$, as desired.
@@
With the notation of Theorem 7.4, the $(R_{p}G_{\bar{0}}b,R_{p}N_{\bar{0}}^{\rho,d}f)$-bimodule $X_{\bar{0}}^{R_{p}}$ and its dual induce a Morita equivalence between $R_{p}G_{\bar{0}}b$ and $R_{p}N_{\bar{0}}^{\rho,d}f$.
Corollary 7.5 answers a question of Kleshchev and Livesey proposed in [22, Remark 8.15] in the highest generality of their question.
@@
Proof. Since there is a finite extension of $\mathbb{Q}_{p}$ containing a primitive $(2n!)$-th root of unity and containing $\sqrt{-1}$, we may assume that $k$ is finite. By [22, Lemma 9.1], the $(\mathcal{O}Gb,\mathcal{O}N^{\rho,d}f)$-bimodule $X^{\mathcal{O}}$ is isomorphic to a direct summand of
@@
for some source idempotents $i\in(\mathcal{O}Gb)^{D}$ and $j\in(\mathcal{O}N^{\rho,d}f)^{D}$, where $U$ is an endopermutation $\mathcal{O}\Delta D$-module. We can regard $U$ as an endopermutation $\mathcal{O}D$-module via the isomorphism $D\cong\Delta D$. By [22, Lemma 9.1], $U$ has an $N_{G}(D)$-stable endosplit $p$-permutation resolution $\mathcal{Y}$; see [30, Section 7] or [24, Definition 7.11.1] for this notion. It follows that $\mathcal{Y}$ is $\mathcal{F}$-stable, where $\mathcal{F}$ is the fusion system on $D$ determined by the source idempotent $i$ of $\mathcal{O}Gb$. By [24, Theorem 9.11.5 (i)], $X^{\mathcal{O}}$ has an endosplit $p$-permutation resolution $Y^{\mathcal{O}}$, and $Y^{\mathcal{O}}$ induces a splendid Rickard equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N^{\rho,d}f$. Let $X^{k}=k\otimes_{\mathcal{O}}X^{\mathcal{O}}$, $Y^{k}=k\otimes_{\mathcal{O}}Y^{\mathcal{O}}$ and $X^{\mathbb{F}_{p}}=\mathbb{F}_{p}\otimes_{\mathbb{Z}_{p}}X^{\mathbb{Z}_{p}}$. By definition, one checks that $Y^{k}$ is an endosplit $p$-permutation resolution of $X^{k}$. By Theorem 7.4, we have $X^{k}\cong k\otimes_{\mathbb{F}_{p}}X^{\mathbb{F}_{p}}$. Then by [28, Theorem 1.3 (b)], there is a splendid complex of $(\mathbb{F}_{p}G\bar{b},\mathbb{F}_{p}N^{\rho,d}\bar{f})$-bimodules $Y^{\mathbb{F}_{p}}$ such that $Y^{k}\cong k\otimes_{\mathbb{F}_{p}}Y^{\mathbb{F}_{p}}$. By [19, Proposition 4.5 (a)], $Y^{\mathbb{F}_{p}}$ induce a splendid Rickard equivalence between $\mathbb{F}_{p}G\bar{b}$ and $\mathbb{F}_{p}N^{\rho,d}\bar{f}$. By lifting property of splendid Rickard equivalences ([30, Theorem 5.2]), $Y^{\mathbb{F}_{p}}$ lifts to a splendid Rickard complex $Y^{\mathbb{Z}_{p}}$ of $(\mathbb{Z}_{p}Gb,\mathbb{Z}_{p}N^{\rho,d}f)$-bimodules.
@@
Let $G=\widetilde{S}_{n}^{+}$ (and hence $G_{\bar{0}}:=\widetilde{A}_{n}$). Then $\mathbb{Z}_{p}G_{\bar{0}}b$ is splendidly Rickard equivalent to $\mathbb{Z}_{p}N^{\rho,d}_{\bar{0}}f$.
@@
where the isomorphism is by using the Mackey formula. Again since $p\nmid|G\times N^{\rho,d}:G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}|$, any vertex of the bimodule $X_{\bar{0}}^{\mathcal{O}}$ has the same order as a vertex of the bimodule $X^{\mathcal{O}}$. It follows that $X_{\bar{0}}^{\mathcal{O}}$ has an endopermutation $)\mathcal{O}({}^{t}(\Delta D$))-source $^{t}U$ for some $t\in G\times N^{\rho,d}/G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}$. Write $t=(t_{1},t_{2})$ with $t_{1}\in G$ and $t_{2}\in N^{\rho,d}$; then $^{t}(\Delta D)=\{({}^{t_{1}}u,{}^{t_{2}}u)\mid u\in D\}$. Since $U$ is $N_{G}(D)$-stable, we see that $^{t_{1}}U$ is $N_{G_{\bar{0}}}({}^{t_{1}}D)$-stable when regarded as an $\mathcal{O}{}^{t_{1}}D$-module. Write $P:={}^{t_{1}}D$, $Q:={}^{t_{2}}D$ and $\varphi:P\cong Q$ the group isomorphism induced by conjugation action of $t_{2}t_{1}^{-1}$. By [24, Proposition 9.7.1], $P$ is a defect group of $\mathcal{O}G_{\bar{0}}b$ and $Q$ is a defect group of $\mathcal{O}N^{\rho,d}_{\bar{0}}f$. Note that $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
@@
Then using the same argument as in [24, page 320], $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
@@
By [22, Theorem 9.7], when $p\equiv 1~{\rm mod}~4$, the $\mathcal{O}D$-source $U$ of $X^{\mathcal{O}}$ is the trivial module. Hence $X^{\mathcal{O}}$ induces a splendid Morita equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N^{\rho,d}d$, and $X^{\mathcal{O}}_{\bar{0}}$ induces a splendid Morita equivalence between $\mathcal{O}G_{\bar{0}}b$ and $\mathcal{O}N^{\rho,d}_{\bar{0}}f$. So by Theorem 7.4 we immediately have the following:
@@
We can now construct splendid Rickard equivalences between $\mathbb{Z}_{p}Gb$ (resp. $\mathbb{Z}_{p}G_{\bar{0}}b$) and its Brauer correspondent in the normaliser of a defect group.
@@
Suppose that $G$ is $\widetilde{S}_{n}^{+}$ or that $p\equiv 1~{\rm mod}~4$ and $G$ is $\widetilde{S}_{n}^{-}$. Let $R_{p}\in\{\mathbb{Z}_{p},\mathbb{F}_{p}\}$. Then $R_{p}Gb$ and $R_{p}G_{\bar{0}}b$ are splendidly Rickard equivalent to their Brauer correspondent in $R_{p}N_{G}(D)$ and $R_{p}N_{G_{0}}(D)$, respectively.
Proof. By the lifting theorem of splendid Rickard equivalences [30, Theorem 5.2], it suffices to consider $R_{p}=\mathbb{F}_{p}$. By Theorem 6.6, there is a splendid Rickard equivalence between the Brauer correspondent of $\mathbb{F}_{p}Gb$ (resp. $\mathbb{F}_{p}G_{\bar{0}}$) in $\mathbb{F}_{p}N_{G}(D)$ (resp. $\mathbb{F}_{p}N_{G_{0}}(D)$) and $\mathbb{F}_{p}N^{\rho,d}f$ (resp. $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$). Now the statements follows from Theorems 7.6, 7.7 and Proposition 4.8.
@@
The first author is supported by National Key R & D Profram of China (No. 2020YFE020 4200) and NSFC (Nos. 12431001, 12350710787). The second author is supported by Fundamental Research Funds for the Central Universities (CCNU24XJ028/25JC025/25JCPT031), and NSFC (Nos. 12501024, 12471016).

### v5

@@ 6 On spin blocks of $\widetilde{S}^{\eta}_{n}$ and $\widetilde{A}_{n}$
The blocks $R_{p}Gf$ and $R_{p}Nf$ are Morita superequivalent. More precisely, there is an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$, together with its dual $(X^{R_{p}})^{*}$, inducing a Morita superequivalence between $R_{p}Gb$ and $R_{p}N^{\rho,d}f$ and fulfilling $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$. The bimodule $X^{R_{p}}$ has the endopermutation $R_{p}\Delta D$-source $U^{R_{p}}:=\Omega_{R_{p}\Delta D_{1}}^{\ell}(R_{p})\otimes_{R_{p}}\cdots\otimes_{R_{p}}\Omega_{R_{p}\Delta D_{d}}^{\ell}(R_{p})$.
Proof. For the first statement, by Proposition 4.5, it suffices to show that there exists an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$. By Remark 6.14 we have an isomorphism of $R_{p}$-superalgebras
$R_{p}N^{\rho,d}f\cong B^{\eta,\rho,0}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}}).$
We let $M^{R_{p}}$ be the $(B^{\eta,\varnothing,1}_{R_{p}},B^{\eta,\varnothing,1}_{R_{p}})$-bimodule
$M^{R_{p}}:=\Omega_{B^{\eta,\varnothing,1}_{R_{p}}\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}})^{\rm sop}}^{\ell}(B^{\eta,\varnothing,1}_{R_{p}}).$
@@
$\begin{split}M_{N^{\rho,d}}^{R_{p}}&:=B^{\eta,\rho,0}_{R_{p}}\boxtimes_{R_{p}}(M^{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})\\ &=B^{\eta,\rho,0}_{R_{p}}\boxtimes_{R_{p}}\Big({\rm Ind}_{(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}},B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})_{S_{d}}}^{(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})\otimes_{R_{p}}(B^{\eta,\varnothing,1}_{R_{p}}\wr_{s}\mathcal{T}^{\eta}_{d,R_{p}})^{\rm sop}}(M^{R_{p}})^{\boxtimes d}_{S_{d}}\Big).\end{split}$
Then by Proposition 3.6, we have $M_{N^{\rho,d}}^{R}\cong R\otimes_{R_{p}}M_{N^{\rho,d}}^{R_{p}}$ as $(RN^{\rho,d}f,RN^{\rho,d}f)$-bisupermodules. Now by Proposition 4.12, there is an $(R_{p}Gb,R_{p}N^{\rho,d}f)$-bisupermodule $X^{R_{p}}$ such that $X^{R}\cong R\otimes_{R_{p}}X^{R_{p}}$, as desired.
Denote by $U^{R_{p}}$ the $R_{p}\Delta D$-module $\Omega_{R_{p}\Delta D_{1}}^{\ell}(R_{p})\otimes_{R_{p}}\cdots\otimes_{R_{p}}\Omega_{R_{p}\Delta D_{d}}^{\ell}(R_{p})$. By Proposition 4.8, we have
$R\otimes_{R_{p}}U^{R_{p}}\cong\Omega_{R\Delta D_{1}}^{\ell}(R)\otimes_{R}\cdots\otimes_{R}\Omega_{R\Delta D_{d}}^{\ell}(R)$
@@
With the notation of Theorem 8.4, the $(R_{p}G_{\bar{0}}b,R_{p}N_{\bar{0}}^{\rho,d}f)$-bimodule $X_{\bar{0}}^{R_{p}}$ induces a Morita equivalence between $R_{p}G_{\bar{0}}b$ and $R_{p}N_{\bar{0}}^{\rho,d}f$ and has an endopermutation source.
Proof. By Theorem 8.4 and [23, Lemma 3.10], $X_{\bar{0}}^{R_{p}}$ induces a Morita equivalence between $R_{p}G_{\bar{0}}b$ and $R_{p}N_{\bar{0}}^{\rho,d}f$. Since ${\rm Res}_{G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}}^{G\times N^{\rho,d}}(X^{R_{p}})=X_{\bar{0}}^{R_{p}}\oplus X^{R_{p}}_{\bar{1}}$, $X^{R_{p}}_{\bar{0}}$ is isomorphic to a direct summand of
${\rm Res}_{G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}}^{G\times N^{\rho,d}}{\rm Ind}_{\Delta D}^{G\times N^{\rho,d}}(U^{R_{p}}).$
[…]
Corollary 8.5 answers a question of Kleshchev and Livesey proposed in [23, Remark 8.15] in the highest generality of their question.
@@
Proof. Since there is a finite extension of $\mathbb{Q}_{p}$ containing a primitive $(2n!)$-th root of unity and containing $\sqrt{-1}$ and $\sqrt{2}$, we may assume that $k$ is finite. By [23, Lemma 9.2], the $(\mathcal{O}Gb,\mathcal{O}N^{\rho,d}f)$-bimodule $X^{\mathcal{O}}$ is isomorphic to a direct summand of
@@
The block $\mathbb{Z}_{p}G_{\bar{0}}b$ is splendidly Rickard equivalent to $\mathbb{Z}_{p}N^{\rho,d}_{\bar{0}}f$.
@@
where the isomorphism is by using the Mackey formula. Again since $p\nmid|G\times N^{\rho,d}:G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}|$, any vertex of the bimodule $X_{\bar{0}}^{\mathcal{O}}$ has the same order with any vertex of the bimodule $X^{\mathcal{O}}$. It follows that $X_{\bar{0}}^{\mathcal{O}}$ has an endopermutation $)\mathcal{O}({}^{t}(\Delta D$))-source $^{t}U$ for some $t\in G\times N^{\rho,d}/G_{\bar{0}}\times N_{\bar{0}}^{\rho,d}$. Write $t=(t_{1},t_{2})$ with $t_{1}\in G$ and $t_{2}\in N^{\rho,d}$; then $^{t}(\Delta D)=\{({}^{t_{1}}u,{}^{t_{2}}u)\mid u\in D\}$. Since $U$ is $N_{G}(D)$-stable, we see that $^{t_{1}}U$ is $N_{G_{\bar{0}}}({}^{t_{1}}D)$-stable when regarded as an $\mathcal{O}{}^{t_{1}}D$-module. Write $P:={}^{t_{1}}D$, $Q:={}^{t_{2}}D$ and $\varphi:P\cong Q$ the group isomorphism induced by conjugation action of $t_{2}t_{1}^{-1}$. By [25, Proposition 9.7.1], $P$ is a defect group of $\mathcal{O}G_{\bar{0}}b$ and $Q$ is a defect group of $\mathcal{O}N^{\rho,d}_{\bar{0}}f$. Note that $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
@@
Then using the same argument as in [25, page 320], $X_{\bar{0}}^{\mathcal{O}}$ is isomorphic to a direct summand of
@@
By [23, Theorem 9.7], when $p\equiv 1~{\rm mod}~4$, the $\mathcal{O}D$-source $U$ of $X^{\mathcal{O}}$ is the trivial module. Hence $X^{\mathcal{O}}$ induces a splendid Morita equivalence between $\mathcal{O}Gb$ and $\mathcal{O}N^{\rho,d}d$, and $X^{\mathcal{O}}_{\bar{0}}$ induces a splendid Morita equivalence between $\mathcal{O}G_{\bar{0}}b$ and $\mathcal{O}N^{\rho,d}_{\bar{0}}f$. So by Theorem 8.4 we immediately have the following:
@@
We can now construct splendid Rickard equivalences between $\mathbb{Z}_{p}Gb$ (resp. $\mathbb{Z}_{p}G_{\bar{0}}b$) and its Brauer correspondent in $\mathbb{Z}_{p}N_{G}(D)$ (resp. $\mathbb{Z}_{p}N_{G_{\bar{0}}}(D)$).
@@
The block $R_{p}Gb$ (resp. $R_{p}G_{\bar{0}}b$) is splendidly Rickard equivalent to its Brauer correspondent in $R_{p}N_{G}(D)$ (resp. $R_{p}N_{G_{0}}(D)$).
Proof. By the lifting theorem of splendid Rickard equivalences (Theorem 5.9), it suffices to consider $R_{p}=\mathbb{F}_{p}$. By Theorem 7.6, there is a splendid Rickard equivalence between the Brauer correspondent of $\mathbb{F}_{p}Gb$ (resp. $\mathbb{F}_{p}G_{\bar{0}}$) in $\mathbb{F}_{p}N_{G}(D)$ (resp. $\mathbb{F}_{p}N_{G_{0}}(D)$) and $\mathbb{F}_{p}N^{\rho,d}f$ (resp. $\mathbb{F}_{p}N^{\rho,d}_{\bar{0}}f$). Now the statements follows from Theorems 8.6, 8.7 and Proposition 5.7.
@@
The first author is supported by National Key R & D Program of China (No. 2020YFE020 4200) and NSFC (Nos. 12431001, 12350710787). The second author is supported by Fundamental Research Funds for the Central Universities (CCNU24XJ028/25JC025/25JCPT031), China Postdoctoral Science Foundation (GZC20252006, 2025T001HB) and NSFC (Nos. 12501024, 12471016).